OpenStudy (anonymous):

factor this here meganomial. 12x³y+21x²y²+16xy³

OpenStudy (anonymous):

xy(12x²+21xy+16y²) so far

OpenStudy (anonymous):

the coefficients also have a common factor of 3

OpenStudy (anonymous):

I'm retarded.

OpenStudy (dumbcow):

meganomial...nice term, haven't heard that used

OpenStudy (angela210793):

16/3???????? O.o

OpenStudy (anonymous):

i will certainly use it. right up there with "trickonometry"

OpenStudy (anonymous):

oh wait, 16doesnotdivideby3...

OpenStudy (anonymous):

oops no i was wrong sorry

OpenStudy (anonymous):

i forgive you satellite :)

OpenStudy (anonymous):

guess you got it right at the start! good work!

OpenStudy (anonymous):

it doesn't simplify further?

OpenStudy (anonymous):

looks like a factored trinomial 0.o a²+ab+b²

OpenStudy (anonymous):

well, you're my last hope myin.

myininaya (myininaya):

\[12(16)=24(8)=6(32)=2(96)=4(48)=\] i don't see we can factor this over the integers but we could factor over the complex

myininaya (myininaya):

would you like me to do this i will do it on paper k?

OpenStudy (anonymous):

KK

OpenStudy (angela210793):

i don't think it can b factorized more than u already have

OpenStudy (anonymous):

hmm, really?

OpenStudy (angela210793):

tht's wht i think....w8 for Myini though

myininaya (myininaya):

ok i have produced a formula to factor things like this

myininaya (myininaya):

\[xy(x+\frac{21y+\sqrt{(21y)^2-4(12)(16y^2)}}{2 \cdot 12})(12x+\frac{21y-\sqrt{(21y)^2-4(12)(16y^2})}{2})\]

myininaya (myininaya):

we can probably simplify this alittle

OpenStudy (anonymous):

holy pellet. I don't think my teacher expects quadratic formula for this lol

myininaya (myininaya):

lol i don't think your teacher wants this either but no you cannot factor this over the integers